OpenStudy (anonymous):

A photon has a wavelength of 6.2 meters. Calculate the energy of the photon in joules. (Planck's constant is 6.626 × 10-34 joule seconds, the speed of light is 2.998 × 108 m/s) REALLY NEED HELP ASAP NEED TO GET THIS UNIT DONE FAST

OpenStudy (shreehari499):

Hey there! @susan12345 ! \(\Huge\tt\color{#CC0000}W\color{#CD0B00}e\color{#CF1700}l\color{#D12200}c\color{#D32E00}o\color{#D53900}m\color{#D74500}e~\color{#DB5C00}t\color{#DD6800}o~\color{#E17F00}O\color{#E38B00}p\color{#E59600}e\color{#E7A200}n\color{#E9AD00}S\color{#EBB900}t\color{#EDC500}u\color{#EFD000}d\color{#F1DC00}y\color{#F3E700}!\color{#F5F300}!\)

OpenStudy (shreehari499):

I will not solve it for you but I will help you out with the formulae and sustitution...

OpenStudy (shreehari499):

We know that E=Hv( where v is the frequency)

OpenStudy (shreehari499):

v=c/Lambda

OpenStudy (shreehari499):

Therefore E= hc/Lambda

OpenStudy (shreehari499):

We know h= Planck's constsnt = 6.626 × 10-34 c = Speed of light = 2.998 × 108 m/s and wavelength Lambda= 6.2m

OpenStudy (shreehari499):

Substitute them to get the answer ...:)